Dan McClure owns a thriving independent bookstore in artsy New Hope, Pennsylvania. He must decide how many copies to order of a new book, Power and Self-Destruction, an exposé on a famous politician’s lurid affairs. Interest in the book will be intense at first and then fizzle quickly as attention turns to other celebrities. The book’s retail price is $20, and the wholesale price is $12. The publisher will buy back the retailer’s leftover copies at a full refund, but McClure Books incurs $4 in shipping and handling costs for each book returned to the publisher. Dan believes his demand forecast can be represented by a normal distribution with a mean of 200 and a standard deviation of 80.

f. Suppose Dan orders 300 copies of the book. What is Dan’s expected leftover inventory? A. 101.64 B. 107.64 C. 114.58 D. 125.68

g. Suppose Dan orders 300 copies of the book. What are Dan’s expected sales? A. 275.69 B. 192.36 C. 258.69 D. 587.25

h. Suppose Dan orders 300 copies of the book. What is Dan’s expected profit? A. $104 B. $1,104 C. $3,140 D. $2,104

i. How many books should Dan order if he wants to achieve a 95 percent in-stock probability? A. 458 B. 758 C. 336 D. 833
Answer #1

f.

Order size, Q = 300

So, z-value = (Q - Mean demand) / Stdev = (300 - 200)/80 = 1.3

The corresponding I(z) from the table = 1.3455

So, expected left-over inventory, I(Q) = I(z)*stdev = 1.3455*80 = 107.64

g.

Expected sales, S(Q) = Q - I(Q) = 300 - 107.64 = 192.36

h.

Cu = Retail price - Wholesale price = 20 - 12 = 8
Co = Shipping cost = 4

Expected profit = Cu * S(Q) - Co * I(Q) = 8*192 - 4*108 = $1,104

i.

F(z) = 0.95 which gives z = 1.7 from the table

So, Q = Mean demand + z * stdev = 200 + 1.7*80 = 336